Person x is 17 person y is 49
How old will they both be when person y is 3 times older than person x.

Answer by stanbon(75887)

Person x is 17 person y is 49
How old will they both be when person y is 3 times older than person x.
--------------
x = 17 ; y = 49
Equation:
49+k = 3(17+k)
49+k = 51+3k
2k = -2
k = -1
-----
x will be 17-1 = 16
y will be 49-1 = 48
------------------
Cheers,
Stan H.
---------------------